OpenStudy (anonymous):

(x + 15) + 2(y + 6). Identify an equivalent expression.

OpenStudy (calculusxy):

First of all, we should try to distribute the 2 to the terms inside of the parenthesis. \[(x + 15) + 2(y + 6) = (x + 15) + 2(y) + 2(6) = x + 15 + 2y + 12\]

OpenStudy (calculusxy):

Now combine all of the like terms ... what do you get afterwards?

OpenStudy (calculusxy):

@rhill

OpenStudy (calculusxy):

combine the like terms in \[x + 15 + 2y + 12 => x + (15 + 12) + 2y\]

OpenStudy (calculusxy):

Just evaluate 15 + 12

OpenStudy (calculusxy):

I think that it would be like: \[\Huge \ x + 27 + 2y\]